# HyDeFuse: Provably Convergent Denoiser-Driven Hyperspectral Fusion

### v1

@@ 3.1 Definitions and Notations
Generally, we use bold lowercase letters ($\boldsymbol{x},\boldsymbol{e}$, etc.) to represent vectors, bold uppercase letters ($\mathbf{X},\mathbf{Z}$, etc.) to represent matrices, and calligraphic letters ($\mathcal{T},\mathcal{W}$, etc.) to denote operators acting on matrices. For any $\boldsymbol{x}\in\mathbb{R}^{n}$, $\operatorname*{diag}(\boldsymbol{x})\in\mathbb{R}^{n\times n}$ denotes a diagonal matrix with $\boldsymbol{x}$ as its diagonal entries. The identity matrix is denoted by $\mathbf{I}$, where the size should be clear from the context. We say that $\mathbf{X}$ is a fixed point of $\mathcal{T}$ if $\mathcal{T}(\mathbf{X})=\mathbf{X}$; the set of fixed points of $\mathcal{T}$ is denoted as $\mathrm{fix}\,(\mathcal{T})$. If $\mathcal{T}$ is linear, $\mathrm{fix}\,(\mathcal{T})$ corresponds to the eigenspace associated with the eigenvalue $1$.
We use $\mathbb{H}=\mathbb{R}^{N_{m}\times L_{s}}$ to denote the space of the matrix-valued variable $\mathbf{X}$ in (3). The reconstruction algorithm will operate in this vector space. To establish convergence, we will work with an inner product in $\mathbb{H}$ and the associated induced norm. In particular, it will be convenient to work with the following inner product on $\mathbb{H}$,
$\big{\langle}\mathbf{X}_{1},\mathbf{X}_{2}\big{\rangle}_{\mathbb{H}}=\mathrm{tr}(\mathbf{X}_{1}^{\top}\mathbf{X}_{2})\qquad(\mathbf{X}_{1},\mathbf{X}_{2}\in\mathbb{H}),$ (4)
where $\mathrm{tr}$ is the trace operator. The columns of $\mathbf{X}\in\mathbb{H}$ are the spectral bands, so the inner product in (4) amounts to computing the standard dot product between the corresponding bands of $\mathbf{X}_{1}$ and $\mathbf{X}_{2}$, followed by summation over all bands. The norm induced by this inner product is given by
@@
We use $\mathcal{L}(\mathbb{H})$ to denote the class of linear operators that map $\mathbb{H}$ into $\mathbb{H}$. We use $\sigma(\mathbf{M})$ and $\sigma(\mathcal{T})$ for the spectrum (set of eigenvalues) of a matrix $\mathbf{M}$ or an operator $\mathcal{T}$. We use $\sigma_{\max}(\mathbf{M})$ for the largest singular value of $\mathbf{M}$. This should not be confused with the notation for the spectrum; the distinction should be clear from the context. We will use the inequality $\|\mathbf{A}\mathbf{B}\|_{\mathbb{H}}\leqslant\sigma_{\max}(\mathbf{A})\|\mathbf{B}\|_{\mathbb{H}}$ in our analysis.
We will use the concept of a self-adjoint operator, a natural generalization of the concept of a symmetric matrix [67]. An operator $\mathcal{T}$ on $(\mathbb{H},\big{\langle}\cdot,\cdot\rangle_{\mathbb{H}})$ is said to be self-adjoint if
$\big{\langle}\mathcal{T}(\mathbf{X}_{1}),\mathbf{X}_{2}\big{\rangle}_{\mathbb{H}}=\big{\langle}\mathbf{X}_{1},\mathcal{T}(\mathbf{X}_{2})\big{\rangle}_{\mathbb{H}}\qquad(\mathbf{X}_{1},\mathbf{X}_{2}\in\mathbb{H}).$
A symmetric matrix or a self-adjoint operator is positive semidefinite if all the eigenvalues are nonnegative. An operator $\mathcal{T}$ on $\mathbb{H}$ is nonexpansive if $\|\mathcal{T}(\mathbf{X}_{1})-\mathcal{T}(\mathbf{X}_{2})\|_{\mathbb{H}}\leqslant\|\mathbf{X}_{1}-\mathbf{X}_{2}\|_{\mathbb{H}}$ for all $\mathbf{X}_{1},\mathbf{X}_{2}\in\mathbb{H}$. In particular, a matrix $\mathbf{M}$ is nonexpansive if $\sigma_{\max}(\mathbf{M})\leqslant 1$. A function $\ell:\mathbb{H}\to\mathbb{R}$ is said to be convex if $\ell(\theta\mathbf{X}_{1}+(1-\theta)\mathbf{X}_{2})\leqslant\theta\ell(\mathbf{X}_{1})+(1-\theta)\ell(\mathbf{X}_{2})$ for all $\theta\in(0,1)$ and $\mathbf{X}_{1},\mathbf{X}_{2}\in\mathbb{H}$.
Figure 2: The processing blocks of HyDeFuse. The hyperspectral image $\mathbf{Y}_{h}$ is upscaled (interpolated) to match the spatial dimension of the multispectral image, which is used to estimate the spectral subspace $\mathbf{E}$. The loss function $\ell$ is derived from the observed HS and MS images and the forward operators $\mathbf{A}$ and $\mathbf{R}$. PnP-PGD is an iterative algorithm where each iteration consists of a gradient step on the loss function followed by regularization using denoiser $\mathcal{D}$. The output of PnP-PGD is $\widehat{\mathbf{X}}$ and $\widehat{\mathbf{Z}}=\widehat{\mathbf{X}}\mathbf{E}$ is the output of HyDeFuse.
3.2 Plug-and-Play Fusion
The PnP framework [42] has its origin in the standard variational framework for solving (3), namely,
@@ 3.2 Plug-and-Play Fusion
where $\ell:\mathbb{H}\to\mathbb{R}$ is the model-based loss function,
@@
and $\varphi:\mathbb{H}\to\mathbb{R}$ is a (convex) regularizer. Since the regularizer $\varphi$ is typically non-smooth, proximal algorithms are used to perform the optimization [65, 8, 68, 69]. For instance, since the loss is differentiable, we can use the proximal gradient descent (PGD) algorithm. Given an initial estimate $\mathbf{X}_{0}\in\mathbb{H}$, the PGD iterations are given by
$\mathbf{X}_{k+1}=\mathrm{prox}_{\gamma\varphi}\big{(}\mathbf{X}_{k}-\gamma\nabla\!\ell(\mathbf{X}_{k})\big{)}\qquad(k=0,1,\ldots),$ (7)
where $\gamma>0$ is the step size, $\nabla\!\ell$ is the gradient of $\ell$, and $\mathrm{prox}_{\gamma\varphi}:\mathbb{H}\to\mathbb{H}$ is the proximal operator of $\gamma\varphi$,
@@
Instead of relying on a classical regularizer and its proximal operator that acts as a Gaussian denoiser [42], the ingenious idea in PnP was to replace the proximal operator in (7) with an arbitrary denoising operator $\mathcal{D}:\mathbb{H}\to\mathbb{H}$. That is, the update in PnP is performed as
$\mathbf{X}_{k+1}=\mathcal{D}\big{(}\mathbf{X}_{k}-\gamma\nabla\!\ell(\mathbf{X}_{k})\big{)}\qquad(k=0,1,\ldots).$ (9)
The point is that denoisers such as BM3D [70], DnCNN [39, 43], and DRUNet [61] are more potent than the proximal operator (8) corresponding to traditional regularizers such as total variation, wavelets, and low-rank [8]. The hope is that a more potent denoiser $\mathcal{D}$ in (9) can produce better reconstructions. However, the challenge lies in the fact that the transition from (7) to (9) is carried out in an ad-hoc manner, with no guarantee that the sequence of iterates $\{\mathbf{X}_{k}\}$ generated by (9) will converge to a meaningful reconstruction. The example in Figure 1 shows that the iterative process can actually fail, leading to divergence. Thus, we need to work with a denoiser $\mathcal{D}$ that not only gives good reconstructions but for which the iterates $\{\mathbf{X}_{k}\}$ are always convergent.
In the rest of the paper, we will refer to (9) as PnP-PGD. We note that this is not a new algorithm and has been used in prior works on PnP [59, 49]. The novelty of the present work is the denoiser we propose next and the proof of convergence of PnP-PGD for this particular denoiser.
We note that the gradient $\nabla\!\ell$ in (7) is with respect to the inner product (4), i.e., the matrix $\nabla\!\ell(\mathbf{X})$ is such that
$\ell(\mathbf{X}+\mathbf{H})=\ell(\mathbf{X})+\langle\nabla\!\ell(\mathbf{X}),\mathbf{H}\big{\rangle}_{\mathbb{H}}+r(\mathbf{H}),$ (10)
where the residual $r:\mathbb{H}\to\mathbb{R}$ is such that $r(\mathbf{H})/\|\mathbf{H}\|_{\mathbb{H}}\to 0$ as $\|\mathbf{H}\|_{\mathbb{H}}\to 0$. Substituting (6) into (10) and performing some calculations, we can verify that
@@
Moreover, since $\mathbf{E}\mathbf{E}^{\top}=\mathbf{I}$, we have
$\nabla\!\ell(\mathbf{X})=\mathbf{A}^{\top}\!\mathbf{A}\mathbf{X}-\mathbf{A}^{\top}\mathbf{Y}_{h}\mathbf{E}^{\top}+\lambda\mathbf{X}(\mathbf{E}\mathbf{R})(\mathbf{E}\mathbf{R})^{\top}-\lambda\mathbf{Y}_{m}(\mathbf{E}\mathbf{R})^{\top}.$ (11)
@@ 3.3 Kernel Denoiser
We now describe the denoiser $\mathcal{D}$ in (9). The noise in hyperspectral images is assumed to be uncorrelated both within and across bands [65], but the bands exhibit strong correlations, with numerous similar patches present within each band. To exploit the correlation across bands, we first denoise the image using a kernel denoiser proposed in [52]. Furthermore, to exploit the presence of similar patches within each band, we process each band of the denoised image separately using another kernel denoiser [42]. Specifically, if we denote the first denoiser as an operator $\mathcal{W}:\mathbb{H}\to\mathbb{H}$ and the second denoiser as $\mathcal{V}:\mathbb{H}\to\mathbb{H}$, then the proposed denoiser $\mathcal{D}$ is given by
@@
We will refer to $\mathcal{W}$ as the high-dimensional kernel denoiser, $\mathcal{V}$ as the bandwise kernel denoiser, and $\mathcal{D}$ as CasKD (cascaded kernel denoiser).
The properties of $\mathcal{W}$ and $\mathcal{V}$ will play a crucial role in the convergence analysis. Therefore, we first give the mathematical description of these denosiers. Both $\mathcal{W}$ and $\mathcal{V}$ are examples of kernel denoisers [71], where each pixel is denoised using a weighted average of its neighboring pixels. The weights are determined using a kernel (or affinity) function, hence the name “kernel denoiser.” The main difference between $\mathcal{W}$ and $\mathcal{V}$ is in the definition of the kernel function.
@@
$\mathbf{K}^{(b)}_{ij}=h(i-j)\phi(\boldsymbol{p}_{i}-\boldsymbol{p}_{j})\qquad(1\leqslant i,j\leqslant N_{m}),$ (13)
@@
  • $\phi$ is a multivariate Gaussian (RBF) kernel function,
  • $h$ is a symmetric hat function supported on a square window around the origin, and
  • $\boldsymbol{p}_{i}$ and $\boldsymbol{p}_{j}$ are the (vectorized) patches around pixels $i$ and $j$ extracted from a guide image, e.g., a surrogate of the ground truth constructed from the observed images [59, 60].
By construction, $\mathbf{K}^{(b)}_{ij}$ takes on large values when pixels $i$ and $j$ and the corresponding patches $\boldsymbol{p}_{i}$ and $\boldsymbol{p}_{j}$ are close. A box function could be used instead of a hat function in (13); the latter guarantees that the kernel matrix is positive semidefinite [42]. In the standard construction of a kernel denoiser [71], the weight matrix is obtained by normalizing $\mathbf{K}^{(b)}$,
@@
where $\boldsymbol{e}$ is the all-ones vector of appropriate length. The problem is that although $\mathbf{K}^{(b)}$ is symmetric, $\mathbf{W}^{(b)}$ may not remain symmetric after normalization –– the product of two symmetric matrices need not be symmetric. As originally proposed in [42], a symmetric weight matrix can be constructed using the following formula, which we continue to denote by the same symbol:
$\mathbf{W}^{(b)}=\frac{1}{\nu}\mathbf{D}^{-\frac{1}{2}}\mathbf{K}^{(b)}\mathbf{D}^{-\frac{1}{2}}+\operatorname*{diag}\Big{(}\boldsymbol{e}-\frac{1}{\nu}\,\hat{\boldsymbol{e}}\Big{)}$, (14)
@@
where $\mathbf{X}_{b}\in\mathbb{R}^{N_{m}}$ and $\mathcal{V}(\mathbf{X})_{b}$ denote the $b$-th band of the input and output of the denoiser.
@@
    Given the input hyperspectral image represented as a matrix $\mathbf{X}$ of size $N_{m}\times L_{s}$, we extract patches around each pixel while considering all the $L_{s}$ bands. Specifically, for each pixel $1\leqslant i\leqslant N_{m}$, we have a patch vector $\boldsymbol{p}_{i}$ of dimension $L_{s}k^{2}$, assuming that a $k\times k$ neighborhood is used around each pixel.
@@
    $\mathbf{K}_{i,j}=h(i-j)\,\sum_{c=1}^{C}\phi(\boldsymbol{p}_{i}-\boldsymbol{\mu}_{c})\phi(\boldsymbol{p}_{j}-\boldsymbol{\mu}_{c}),\qquad(1\leqslant i,j\leqslant N_{m}),$ (16)
    where $\phi$ is again a Gaussian kernel. Note that $\mathbf{K}_{i,j}$ assumes relatively large values when $\boldsymbol{p}_{i}$ and $\boldsymbol{p}_{j}$ belong to the same cluster.
@@
    $\mathbf{W}=\frac{1}{\nu}\mathbf{D}^{-\frac{1}{2}}\mathbf{K}\mathbf{D}^{-\frac{1}{2}}+\operatorname*{diag}\Big{(}\boldsymbol{e}-\frac{1}{\nu}\,\hat{\boldsymbol{e}}\Big{)}$, (17)
@@
The primary motivation behind the construction in (16) is that it facilitates the development of a fast convolution-based algorithm for computing (18). Details on this and other implementation aspects can be found in [52]. Rather, what is more important is the following observation.
@@
The matrices $\mathbf{W}^{(b)}$ and $\mathbf{W}$ in (14) and (17) are symmetric positive semidefinite, stochastic, and irreducible.
Stochastic means $\mathbf{W}_{ij}$ are nonnegative, and all the rows sum to one. By irreducible, we mean that for all $i,j$, there exists $k\geqslant 1$, such that $(\mathbf{W}^{k})_{ij}>0$.
@@
The technical details behind Proposition 1 can be found in [42, Sec. IV]. However, one can directly verify from (17) that $\mathbf{W}$ is symmetric, nonnegative, and $\mathbf{W}\boldsymbol{e}=\boldsymbol{e}$ (stochastic). It is also not difficult to show that $\mathbf{W}$ –– for that matter, any operator that performs local averaging on the signal using strictly positive weights –– is irreducible. The only tricky property is the positive definite property; this can be deduced using the Fourier transform and Bochner’s theorem [42].
We note that different orderings and combinations of $\mathcal{V}$ and $\mathcal{W}$, such as $\mathcal{W}\circ\mathcal{V}$, could have been used in (12). However, we empirically found that $\mathcal{V}\circ\mathcal{W}$ gives slightly better results. In either case, the denoising performance surpasses that of using $\mathcal{V}$ or $\mathcal{W}$ alone. This is demonstrated in Table 1 and Figure 3. Moreover, since ultimately we wish to use CasKD to perform fusion (and not for standalone denoising), we have compared its regularization capacity for HyDeFuse with the component denoisers in Table 9. We see a significant gap in the fusion quality.
@@
20/255 $\mathcal{W}\circ\mathcal{V}$ 38.01 0.986 36.87 0.982
@@
CasKD $\mathbf{38.06}$ $\mathbf{0.987}$ $\mathbf{37.06}$ $\mathbf{0.982}$
@@
50/255 $\mathcal{W}\circ\mathcal{V}$ 32.38 0.955 30.73 0.915

### v2

@@ 2.4 Definitions and Notations
We use bold lowercase letters ($\boldsymbol{x}$, $\boldsymbol{e}$, etc.) to denote vectors, bold uppercase letters (e.g., $\mathbf{X}$, $\mathbf{Z}$) to denote matrices, and calligraphic letters ($\mathcal{T}$, $\mathcal{W}$, etc.) to denote operators acting on matrices. For any $\boldsymbol{x}\in\mathbb{R}^{n}$, $\operatorname*{diag}(\boldsymbol{x})\in\mathbb{R}^{n\times n}$ denotes the diagonal matrix with $\boldsymbol{x}$ on its diagonal. The identity matrix is denoted by $\mathbf{I}$, with its dimension understood from the context.
We use $\mathbb{H}=\mathbb{R}^{N_{m}\times L_{s}}$ to denote the space of the matrix-valued variable $\mathbf{X}$ in (3). The reconstruction algorithm operates in this vector space. To analyze convergence, we equip $\mathbb{H}$ with an inner product and its induced norm. Specifically, we use the inner product
$\langle\mathbf{X}_{1},\mathbf{X}_{2}\rangle_{\mathbb{H}}=\mathrm{tr}(\mathbf{X}_{1}^{\top}\mathbf{X}_{2}),\qquad\mathbf{X}_{1},\mathbf{X}_{2}\in\mathbb{H},$ (4)
where $\mathrm{tr}$ denotes the trace operator. The norm induced by this inner product is given by
@@
Since the columns of $\mathbf{X}\in\mathbb{H}$ represent spectral bands, the inner product (4) amounts to computing the standard dot product between the corresponding bands of $\mathbf{X}_{1}$ and $\mathbf{X}_{2}$, followed by summation over all bands. In particular,
$\|\mathbf{X}\|^{2}_{\mathbb{H}}=\sum_{i,j}|\mathbf{X}_{ij}|^{2},$
which is just the squared Frobenius norm of $\mathbf{X}$.
We use $\mathcal{L}(\mathbb{H})$ to denote the class of linear operators that map $\mathbb{H}$ into $\mathbb{H}$. We use $\sigma(\mathbf{M})$ and $\sigma(\mathcal{T})$ for the spectrum (set of eigenvalues) of a matrix $\mathbf{M}$ or an operator $\mathcal{T}$. We use $\sigma_{\max}(\mathbf{M})$ for the largest singular value of $\mathbf{M}$.
We say that $\mathbf{X}$ is a fixed point of an operator $\mathcal{T}$ if $\mathcal{T}(\mathbf{X})=\mathbf{X}$. The set of all fixed points of $\mathcal{T}$ is denoted by $\mathrm{fix}\,(\mathcal{T})$. If $\mathcal{T}$ is linear, then $\mathrm{fix}\,(\mathcal{T})$ coincides with the eigenspace corresponding to the eigenvalue $1$.
We will need the concept of a self-adjoint operator, which generalizes the idea of a symmetric matrix [70]. An operator $\mathcal{T}$ on $(\mathbb{H},\langle\cdot,\cdot\rangle_{\mathbb{H}})$ is said to be self-adjoint if
$\langle\mathcal{T}(\mathbf{X}_{1}),\mathbf{X}_{2}\rangle_{\mathbb{H}}=\langle\mathbf{X}_{1},\mathcal{T}(\mathbf{X}_{2})\rangle_{\mathbb{H}},\qquad\mathbf{X}_{1},\mathbf{X}_{2}\in\mathbb{H}.$
A self-adjoint operator (or a symmetric matrix) is called positive semidefinite if all its eigenvalues are nonnegative.
An operator $\mathcal{T}:\mathbb{H}\to\mathbb{H}$ is said to be $L$-Lipschitz if there exists $L>0$ such that
$\|\mathcal{T}(\mathbf{X}_{1})-\mathcal{T}(\mathbf{X}_{2})\|_{\mathbb{H}}\leqslant L\|\mathbf{X}_{1}-\mathbf{X}_{2}\|_{\mathbb{H}},\qquad(\mathbf{X}_{1},\mathbf{X}_{2}\in\mathbb{H}).$
In particular, $\mathcal{T}$ is a contraction operator if $L<1$, and a nonexpansive operator if $L=1$. A matrix $\mathbf{M}$ is nonexpansive if and only if $\sigma_{\max}(\mathbf{M})\leqslant 1$.
A function $\ell:\mathbb{H}\to\mathbb{R}$ is said to be convex if $\ell(\theta\mathbf{X}_{1}+(1-\theta)\mathbf{X}_{2})\leqslant\theta\ell(\mathbf{X}_{1})+(1-\theta)\ell(\mathbf{X}_{2})$ for all $\theta\in(0,1)$ and $\mathbf{X}_{1},\mathbf{X}_{2}\in\mathbb{H}$. Moreover, $\ell$ is said to be $\beta$-smooth if it is differentiable and its gradient operator $\nabla\!\ell:\mathbb{H}\to\mathbb{H}$ is $\beta$-Lipschitz.
A matrix $\mathbf{W}$ is called stochastic if all its entries are nonnegative and each row sums to $1$. It is said to be irreducible if, for every pair of indices $i$ and $j$, there exists $k\geqslant 1$ such that $(\mathbf{W}^{k})_{ij}>0$. Equivalently, the directed graph associated with $\mathbf{W}$ is strongly connected [70].
3 Fusion Algorithm
3.1 Plug-and-Play Framework
We begin by introducing the Plug-and-Play (PnP) framework, which forms the foundation of the proposed fusion algorithm. The pipeline of our method is illustrated in Figure 2. The design and properties of the kernel-based denoiser $\mathcal{D}$ are discussed in the following subsection.
The PnP framework [42] originates from the classical optimization formulation for solving (3):
@@ 3.1 Plug-and-Play Framework
where the model-based loss function $\ell:\mathbb{H}\to\mathbb{R}$ is given by
@@
and $\varphi:\mathbb{H}\to\mathbb{R}$ is a (convex) regularizer. Since $\varphi$ is typically non-smooth, proximal algorithms are used to perform the optimization [68, 8, 71, 72]. The loss $\ell$ is convex, as it is the sum of convex quadratic functions.
Since $\ell$ is differentiable, the proximal gradient descent (PGD) algorithm can be used to solve the optimization problem (5). For a given initialization $\mathbf{X}_{0}\in\mathbb{H}$, the PGD iterations are given by
$\mathbf{X}_{k+1}=\mathrm{prox}_{\gamma\varphi}\big(\mathbf{X}_{k}-\gamma\nabla\!\ell(\mathbf{X}_{k})\big)\qquad(k\geqslant 0),$ (7)
where $\gamma>0$ is the step size, $\nabla\!\ell$ is the gradient of $\ell$, and $\mathrm{prox}_{\gamma\varphi}:\mathbb{H}\to\mathbb{H}$ is the proximal operator of $\gamma\varphi$ given by
@@
Figure 2: The processing pipeline of HyDeFuse, the proposed fusion algorithm. The HS image $\mathbf{Y}_{h}$ is first upscaled (interpolated) to match the spatial resolution of the MS image, which is then used to estimate the spectral subspace $\mathbf{E}$. The loss function $\ell$ is formulated using $\mathbf{E}$, the observed HS and MS images ($\mathbf{Y}_{h}$ and $\mathbf{Y}_{m}$), and the forward operators $\mathbf{A}$ and $\mathbf{R}$. The reconstruction is performed using the iterative PnP-PGD algorithm, where each iteration alternates between a gradient step on $\ell$ and regularization via the proposed hyperspectral denoiser $\mathcal{D}$. The output of PnP-PGD is $\widehat{\mathbf{X}}$ and $\widehat{\mathbf{Z}}=\widehat{\mathbf{X}}\mathbf{E}$ is the output of HyDeFuse.
Instead of relying on a classical regularizer and its proximal operator, which typically acts as a Gaussian denoiser [42], the key innovation in the Plug-and-Play (PnP) framework is to replace the proximal operator in (7) with a generic denoising operator $\mathcal{D}:\mathbb{H}\to\mathbb{H}$. Accordingly, the PnP update is given by
$\mathbf{X}_{k+1}=\mathcal{D}\big(\mathbf{X}_{k}-\gamma\nabla\ell(\mathbf{X}_{k})\big),\qquad k=0,1,\ldots$ (9)
The motivation behind this substitution is that denoisers such as DnCNN [39, 43], and DRUNet [64] are often far more effective than the proximal operators in (8) derived from classical regularizers such as total variation, wavelet sparsity, or low-rank models [8]. The expectation is that employing a stronger denoiser $\mathcal{D}$ in (9) leads to higher-quality reconstructions.
The main technical challenge, however, is that the transition from (7) to (9) is not theoretically justified, with no convergence guarantee on the sequence $\{\mathbf{X}_{k}\}$ produced by (9). Indeed, as illustrated in Figure 1, the iterative process can sometimes diverge. Hence, it is essential to design a denoiser $\mathcal{D}$ that not only performs well empirically but also promotes convergence of the iterative scheme.
In the rest of the paper, we refer to (9) as PnP-PGD. This algorithm is not new and has been used in previous PnP studies [54, 62]. The novelty of our work lies in the new denoiser we introduce next and in establishing the convergence of PnP-PGD when this denoiser is used.
We now derive the expression for the gradient of the loss function $\ell$ with respect to the inner product defined in (4). Specifically, its gradient $\nabla\!\ell(\mathbf{X})$ is defined as the unique element in $\mathbb{H}$ satisfying
$\ell(\mathbf{X}+\mathbf{H})=\ell(\mathbf{X})+\langle\nabla\!\ell(\mathbf{X}),\mathbf{H}\rangle_{\mathbb{H}}+r(\mathbf{H}),$ (10)
where the residual $r:\mathbb{H}\to\mathbb{R}$ is such that $r(\mathbf{H})/\|\mathbf{H}\|_{\mathbb{H}}\to 0$ as $\|\mathbf{H}\|_{\mathbb{H}}\to 0$. Substituting (6) into (10) and performing straightforward calculations, we obtain
@@
Since $\mathbf{E}\mathbf{E}^{\top}=\mathbf{I}$, this expression can be written more compactly as
$\nabla\!\ell(\mathbf{X})=\mathcal{K}(\mathbf{X})-\mathbf{A}^{\top}\mathbf{Y}_{h}\mathbf{E}^{\top}-\lambda\mathbf{Y}_{m}(\mathbf{E}\mathbf{R})^{\top},$ (11)
where $\mathcal{K}\in\mathcal{L}(\mathbb{H})$ is defined as
$\mathcal{K}(\mathbf{X})=\mathbf{P}_{1}\mathbf{X}+\mathbf{X}\mathbf{P}_{2},\qquad\mathbf{P}_{1}:=\mathbf{A}^{\top}\mathbf{A},\ \mathbf{P}_{2}:=\lambda(\mathbf{E}\mathbf{R})(\mathbf{E}\mathbf{R})^{\top}.$ (12)
In other words, $\nabla\!\ell$ is an affine operator. This will play a crucial role in our subsequent analysis.
@@ 3.2 Kernel Denoiser
We now describe the denoiser $\mathcal{D}$ in (9). The noise in hyperspectral images is generally assumed to be uncorrelated both within and across bands [68]. However, the spectral bands themselves exhibit strong correlations, and numerous similar patches occur within each band. To exploit the inter-band correlations, we first denoise the image using the kernel-based denoiser proposed in [47]. To further leverage the intra-band similarity, each band of the denoised image is then processed separately using another kernel denoiser [42].
Specifically, let $\mathcal{W}:\mathbb{H}\to\mathbb{H}$ denote the first (high-dimensional) denoiser and $\mathcal{V}:\mathbb{H}\to\mathbb{H}$ denote the second (bandwise) denoiser. The overall denoising operator $\mathcal{D}$ is defined as
@@
We refer to $\mathcal{W}$ as the high-dimensional kernel denoiser, $\mathcal{V}$ as the bandwise kernel denoiser, and $\mathcal{D}$ as CasKD (Cascaded Kernel Denoiser).
The mathematical properties of $\mathcal{W}$ and $\mathcal{V}$ are central to our convergence analysis. We therefore begin by describing these denoisers in detail. Both $\mathcal{W}$ and $\mathcal{V}$ are kernel denoisers [73], in which each pixel is denoised by taking a weighted average of its neighboring pixels. The weights are computed using a kernel (or affinity) function. The key difference between $\mathcal{W}$ and $\mathcal{V}$ lies in the definition of their respective kernel functions.
@@
$\mathbf{K}^{(b)}_{ij}=h(i-j)\phi_{2}(\boldsymbol{p}_{i}-\boldsymbol{p}_{j})\qquad(1\leqslant i,j\leqslant N_{m}),$ (14)
@@
  • $\boldsymbol{p}_{i}$ and $\boldsymbol{p}_{j}$ are the (vectorized) patches around pixels $i$ and $j$ that are extracted from a guide image, a surrogate of the ground truth constructed from the observed images [54, 63].
  • $\phi_{2}$ is a multivariate Gaussian (RBF) kernel,
    $\phi_{2}(p_{i},p_{j})=\exp\left(-\frac{\|p_{i}-p_{j}\|_{2}^{2}}{2\sigma_{2}^{2}}\right),$
    with $\sigma_{2}$ denoting the standard deviation and $\|\cdot\|_{2}$ representing the $\ell_{2}$ norm on the patch space.
  • $h$ is a symmetric hat function supported on a square window around the origin.
By construction, $\mathbf{K}^{(b)}_{ij}$ takes on large values when pixels $i$ and $j$ and the corresponding patches $\boldsymbol{p}_{i}$ and $\boldsymbol{p}_{j}$ are similar. A simple box function could be used instead of a hat function in (14); however, the latter guarantees that the kernel matrix is positive semidefinite [42].
In the standard construction of a kernel denoiser [73], the weight matrix is obtained by normalizing $\mathbf{K}^{(b)}$,
@@
where $\boldsymbol{e}$ is the all-ones vector of appropriate length. The problem is that although $\mathbf{K}^{(b)}$ is symmetric, $\mathbf{W}^{(b)}$ may not remain symmetric after normalization, the product of two symmetric matrices need not be symmetric. As originally proposed in [42], a symmetric weight matrix can be constructed using the following formula, which we continue to denote by the same symbol:
[…]
@@
where $\mathbf{X}_{b}\in\mathbb{R}^{N_{m}}$ and $\mathcal{V}(\mathbf{X})_{b}$ denote the $b$-th band of the input and the output of the denoiser.
@@
    Given the input hyperspectral image represented as a matrix $\mathbf{X}$ of size $N_{m}\times L_{s}$, we extract patches around each pixel while considering all the $L_{s}$ bands. Specifically, for each pixel $1\leqslant i\leqslant N_{m}$, we have a patch vector $\boldsymbol{p}_{i}$ of dimension $k^{2}L_{s}$, assuming that a $k\times k$ neighborhood is used around each pixel.
@@
    $\mathbf{K}_{ij}=h(i-j)\,\sum_{c=1}^{C}\phi_{1}(\boldsymbol{p}_{i}-\boldsymbol{\mu}_{c})\phi_{1}(\boldsymbol{p}_{j}-\boldsymbol{\mu}_{c}),\qquad(1\leqslant i,j\leqslant N_{m}),$ (17)
    where $\phi_{1}$ is again a Gaussian kernel,
    $\phi_{1}(p_{i},\mu_{c})=\exp\left(-\frac{\|p_{i}-\mu_{c}\|_{2}^{2}}{2\sigma_{1}^{2}}\right),$
    with $\sigma_{1}$ denoting the standard deviation of $\phi_{1}$.
    By construction, $\mathbf{K}_{ij}$ assumes relatively large values when $\boldsymbol{p}_{i}$ and $\boldsymbol{p}_{j}$ belong to the same cluster.
@@
[…]
@@
The main motivation behind the construction in (17) is that it allows a fast, convolution-based computation of (19). Further implementation details are provided in [47]. What is more important, however, is the following observation.
@@
The matrices $\mathbf{W}^{(b)}$ and $\mathbf{W}$ in (15) and (18) are (symmetric) positive semidefinite, stochastic, and irreducible.
@@
The technical details behind Proposition 1 can be found in [42, Sec. IV]. However, one can directly verify from (18) that $\mathbf{W}$ is symmetric, nonnegative, and $\mathbf{W}\boldsymbol{e}=\boldsymbol{e}$ (stochastic). It is also not difficult to show that $\mathbf{W}$ is irreducible; in fact, any operator that performs local averaging on the signal using strictly positive weights is irreducible. The only tricky property is the positive definite property; this can be deduced using the positivity of the Fourier transform of a Gaussian and Bochner’s theorem [42].
We remark that a different ordering of $\mathcal{V}$ and $\mathcal{W}$, such as $\mathcal{W}\circ\mathcal{V}$, can also be used in (13). In practice, different orderings produce similar performance for both denoising and fusion tasks, and our theoretical results hold for all such combinations. For consistency, we use a single ordering throughout this work. In general, the cascaded versions outperform using $\mathcal{V}$ or $\mathcal{W}$ alone, as shown in Table 1 and Figure 3.
@@
20/255 $\mathcal{W}\circ\mathcal{V}$ 38.01 0.986 36.87 $\mathbf{0.982}$
@@
$\mathcal{V}\circ\mathcal{W}$ (CasKD) $\mathbf{38.06}$ $\mathbf{0.987}$ $\mathbf{37.06}$ $\mathbf{0.982}$
@@
50/255 $\mathcal{W}\circ\mathcal{V}$ 32.38 $\mathbf{0.955}$ 30.73 0.915